# Monitoring Risks in Test-Time Adaptation

### v1

@@ A.4 Ablation on Calibration Set Size
Future work may explore reducing computational overhead by performing calibration less frequently—evaluating only at fixed intervals or adaptively triggering calibration after an observed increase in risk, rather than at every time point
@@
A.5 Extended Comparison to Baselines
@@
Fig. 8 displays the estimated test risk across datasets and TTA methods. As expected, the oracle, supervised lower bound on the test risk, $\hat{L}_{t}^{a}$ ( ), reliably flags risk violations without causing false alarms across all datasets and TTA methods. In contrast, the naive plug-in bound $\hat{L}_{t}^{c}$ ( ) triggers a false alarm on the in-distribution ImageNet test stream for CoTTA, despite the test risk remaining below the alarm threshold. This is unsurprising, as $\hat{L}_{t}^{c}$ lacks formal guarantees on the false alarm rate. While it yields reasonable risk estimates for the other TTA methods on ImageNet ID, as well as on ImageNet-C severity 5 and FMoW-Time, it fails to detect risk violations on Yearbook across all TTA methods. Even though not originally proposed for TTA, we extend the unsupervised test risk lower bound by Amoukou et al. [1], $\hat{L}_{t}^{d}$ to the TTA setting to enable comparison on this plot. We note that it behaves poorly with TTA methods. $\hat{L}_{t}^{d}$ ( ), also triggers a false alarm on ImageNet ID for CoTTA. For other TTA methods, it is largely unresponsive resulting in a consistently loose lower bound on the estimated true test risk $\hat{\bar{R}}_{t}$ ( ). This looseness leads to missed alarms on the severe shift of ImageNet-C severity 5 for 3 out of 5 TTA methods. Furthermore, we observe that the required assumption of their method ( ) is violated in nearly every practical setting.
@@ B.1 Choice of Confidence Sequences
For an introduction to confidence sequences, we refer the interested reader to Howard et al. [13] and Appendix E of Podkopaev and Ramdas [30]. Throughout this paper, we use a Hoeffding confidence interval to estimate the upper bound on the source risk. Accordingly, the finite-sample penalty term is given by $w_{0}=\sqrt{\log{(1/\alpha_{\text{source}})}/N_{\text{cal}}}$. For the lower bound on the test risk, we use the conjugate-mixture empirical Bernstein (CM-EB) confidence sequence proposed in Theorem 4 of Howard et al. [13], chosen for its minimal assumptions. To obtain the finite-sample terms $w_{t}$ for $t\geq 1$—which depend on $\alpha_{\text{test}}$ and the empirical variance of the observed sequence—we use the gamma-exponential mixture bound from Proposition 9 in Howard et al. [13].
@@ B.2 Unsupervised Lower Bound Derivation (Propositon 1)
The proof technique is inspired by the derivation presented in Amoukou et al. [1]; see Eqs. (12)–(15) in their paper. To derive a lower bound on the true running test risk, we first apply Markov’s inequality and then invoke Assumption 1:
@@ B.3 PFA Control Guarantee
The proof closely follows the PFA proof for the supervised alarm from Podkopaev and Ramdas [30], see Appendix D there. To show the PFA guarantee we proceed as: [^3] To simplify notation, we omit all arguments of the relevant risks and confidence sequences in the proof (e.g., we abbreviate $U({\mathbf{z}}_{0})$ as $U$).
@@ D.1 Datasets
  • ImageNet-C [12]: This dataset applies 15 types of algorithmic corruptions (e.g., Gaussian noise, blur, weather effects, digital distortions) at five severity levels to the original ImageNet [59] validation set. The dataset preserves the original 1,000-class classification task, using the same labels and image resolutions. In our setup, we focus on Gaussian noise corruption.
  • Yearbook [8]: This dataset contains portraits of American high school students taken over eight decades, capturing changes in visual appearance due to evolving beauty standards, cultural norms, and demographics. We follow the Wild-Time preprocessing and evaluation protocol [45], resulting in 33,431 grayscale images (32×32 pixels) labeled with binary gender. Images from 1930–1969 are used for training, and those from 1970–2013 for testing.
  • FMoW-Time: The Functional Map of the World (FMoW) dataset [60] consists of 224×224 RGB satellite images categorized into 62 land-use classes. Distribution shift arises from technological and economic changes that alter land usage over time. FMoW-Time [45] is a temporal split of FMoW-WILDS [60, 61], dividing 141,696 images into a training period (2002–2014) and a testing period (2015–2017).
@@ D.2 TTA Methods
  • TENT[42] updates normalization layers by minimizing test entropy.
  • SHOT [19] adapts normalization layers using information maximization and self-supervised pseudo-labeling to align target representations with a frozen source classifier.
  • SAR [27] updates normalization layers via an entropy minimization objective. It filters out high-entropy samples and guides adaptation toward flatter minima.
  • CoTTA [41] updates all model parameters using a student-teacher approach on augmentation averaged predictions. It also employs stochastic weight restoration to mitigate forgetting.
  • T3A [15] adjusts only the final linear classifier by computing class-wise pseudo-prototypes from confident, normalized representations.
  • STAD [35] updates only the last linear layer by tracking the evolution of feature representations with a probabilistic state-space model.
@@ D.3 Implementation Details
We use the confseq package [62] by [13] to compute the conjugate-mixture empirical Bernstein confidence lower bound on the target risk. This confidence sequence framework supports tuning for an intrinsic time $t_{\text{opt}}$, which we set by default to the first 25% of the sequence length for all experiments. To implement the baseline $\hat{L}_{t}^{d}$ from Amoukou et al. [1], we use the same loss proxy—uncertainty—as in our method.
@@ Appendix E Further Related Work
aims to assess model performance on unlabeled test data, which is often subject to distribution shift [63, 64]. This is typically achieved via model uncertainty [65, 4, 66, 67, 68] or model disagreement [69, 70, 16, 33, 18, 71, 72]. Uncertainty-based methods exploit the predictive distribution of the model— for example through the maximum class probability [65] or the true class probability [4]—and learn a threshold to distinguish correctly from incorrectly predicted samples [66, 67, 68]. Disagreement-based error prediction methods leverage the theoretical equivalence between model disagreement and test error under calibration [69, 16]. However, these methods often require training multiple models—sometimes even from different architectures [70, 69]. Closest to our work [18, 31, 73], estimate the accuracy of TTA methods based on disagreement. Notably, by exploiting theoretical results from [69], Lee et al. [18] proposes an accuracy estimation method based on dropout disagreement. They differ from our work by (i) providing an estimator of test risk directly while we are interested in signaling a significant increase in test risk compared to the source risk; as such (ii) their method does not come with guarantees on the false alarm rate; and (iii) they require calibration (their Definition 3.3) to preserve theoretical validity of their risk estimator while we rely on separability of high and low error samples (Assumption 1).
@@
Recent work has identified several scenarios where TTA methods tend to degrade. These include adaptation under non-stationary test distributions [41, 74, 46, 53, 6, 22, 75, 76, 35], label shift [10, 26, 3, 46], mixed domains within a test batch [27, 53, 22, 76], small test batch sizes [27, 53, 22, 6, 35], and adaptation in the presence of malicious samples [29, 43]. Most such work on TTA robustness has focused on proposing more robust TTA methods and developing evaluation benchmarks [10, 22, 76, 46, 75]. Liu et al. [77] analyze failure cases of the related test-time training paradigm, which requires a self-supervised objective during training. They derive an upper bound on test risk dependent on the effectiveness of the self-supervised loss. In contrast, our approach does not require any modification to the training procedure and provides guarantees that hold regardless of the model’s original training objective. Also related to our work is research on TTA model collapse—where models degenerate to trivial solutions during adaptation [31, 27, 18, 78]—and efforts to identify optimal reset mechanisms that revert the model back to its source parameters during deployment [26, 18, 78]. In contrast, we propose a general-purpose monitoring tool that provides statistical guarantees on risk control for arbitrary TTA methods. Rather than focusing on a specific mitigation strategy, our tool can inform a range of interventions—such as resetting the model to its source for continued adaptation or taking it offline entirely for retraining [79].

### v2

@@ A.3 Ablation on Calibration Set Size and Calibration Frequency
Another way to reduce the computational overhead of the monitoring tool is to decrease the frequency of the online calibration procedure described in § 3.4. Instead of evaluating the adapted model on the calibration set and selecting a new threshold $\lambda_{k}$ after every adaptation step $k$, one can perform this threshold selection only periodically, reusing the most recently calibrated threshold for the intermediate monitoring intervals. This reduces the number of calibration evaluations while maintaining continuous monitoring of the model’s performance.
Figure 8: Estimated test risk of TENT on ImageNet-C (ID to Severity 5) when performing the online calibration step every 1,2,5,10,20{1,2,5,10,20} adaptation steps. Less frequent calibration results in only a small increase in detection delay, indicating a good trade-off between detection delay and computational overhead.
To assess the impact of such reduced calibration frequency, we monitor TENT performance on the ImageNet-C test stream with different intervals—performing calibration every 1,2,5,10,20{1,2,5,10,20} adaptation steps. Fig. 8 shows that less frequent calibration leads to only a small increase in detection delay. This suggests that our monitoring framework remains effective even when calibration is performed at coarser temporal resolutions. Future work may further explore adaptive calibration strategies that trigger re-calibration only when a significant increase in estimated risk is detected, rather than at fixed time intervals.
A.4 Extended Comparison to Baselines
@@
Fig. 9 displays the estimated test risk across datasets and TTA methods. As expected, the oracle, supervised lower bound on the test risk, $\hat{L}_{t}^{a}$ ( ), reliably flags risk violations without causing false alarms across all datasets and TTA methods. In contrast, the naive plug-in bound $\hat{L}_{t}^{c}$ ( ) triggers a false alarm on the in-distribution ImageNet test stream for CoTTA, despite the test risk remaining below the alarm threshold. This is unsurprising, as $\hat{L}_{t}^{c}$ lacks formal guarantees on the false alarm rate. While it yields reasonable risk estimates for the other TTA methods on ImageNet ID, as well as on ImageNet-C severity 5 and FMoW-Time, it fails to detect risk violations on Yearbook across all TTA methods. Even though not originally proposed for TTA, we extend the unsupervised test risk lower bound by jpmorgan2024sequential, $\hat{L}_{t}^{d}$ to the TTA setting to enable comparison on this plot. We note that it behaves poorly with TTA methods. $\hat{L}_{t}^{d}$ ( ), also triggers a false alarm on ImageNet ID for CoTTA. For other TTA methods, it is largely unresponsive resulting in a consistently loose lower bound on the estimated true test risk $\hat{\bar{R}}_{t}$ ( ). This looseness leads to missed alarms on the severe shift of ImageNet-C severity 5 for 3 out of 5 TTA methods. Furthermore, we observe that the required assumption of their method ( ) is violated in nearly every practical setting.
@@
A.5 Results under Label Shift
As our theoretical framework does not make any assumptions about the nature of the distribution shift (see § 2), it naturally extends to scenarios with shifting label distributions. To explicitly validate this, we conducted additional experiments on the Yearbook dataset, where we induced controlled label shift by reordering test samples according to their class labels, following the setup of [gong2022note, schirmer2024test].
Fig. 10 compares our monitoring results for Yearbook with (right) and without (left) label shift. Test-time adaptation methods perform worse under label shift, as reflected by the higher empirical risk $\hat{\bar{R}}_{t}$ ( ). Our method reliably detects this deterioration, triggering an earlier alarm across all TTA methods. While Assumption 1 becomes slightly looser in this setting, the monitor still raises an alert, demonstrating that our framework effectively captures failures of TTA methods caused by label shift.
Figure 10: Estimated test risk on Yearbook. We compare a test stream affected only by covariate shift (left) with one that additionally exhibits label shift (right). Under label shift, TTA methods show higher empirical risk, while our risk monitor detects the degradation earlier and raises an alarm accordingly.
@@ B.1 Choice of Confidence Sequences
For an introduction to confidence sequences, we refer the interested reader to howard2021time and Appendix E of podkopaev2021tracking. Throughout this paper, we use a Hoeffding confidence interval to estimate the upper bound on the source risk. Accordingly, the finite-sample penalty term is given by $w_{0}=\sqrt{\log{(1/\alpha_{\text{source}})}/N_{\text{cal}}}$. For the lower bound on the test risk, we use the conjugate-mixture empirical Bernstein (CM-EB) confidence sequence proposed in Theorem 4 of howard2021time, chosen for its minimal assumptions. To obtain the finite-sample terms $w_{t}$ for $t\geq 1$—which depend on $\alpha_{\text{test}}$ and the empirical variance of the observed sequence—we use the gamma-exponential mixture bound from Proposition 9 in howard2021time.
@@ B.2 Unsupervised Lower Bound Derivation (Propositon 1)
The proof technique is inspired by the derivation presented in jpmorgan2024sequential; see Eqs. (12)–(15) in their paper. To derive a lower bound on the true running test risk, we first apply Markov’s inequality and then invoke Assumption 1:
@@ B.3 PFA Control Guarantee
The proof closely follows the PFA proof for the supervised alarm from podkopaev2021tracking, see Appendix D there. To show the PFA guarantee we proceed as: [^3] To simplify notation, we omit all arguments of the relevant risks and confidence sequences in the proof (e.g., we abbreviate $U({\mathbf{z}}_{0})$ as $U$).
@@ D.1 Datasets
  • ImageNet-C [hendrycks2019benchmarking]: This dataset applies 15 types of algorithmic corruptions (e.g., Gaussian noise, blur, weather effects, digital distortions) at five severity levels to the original ImageNet [deng2009imagenet] validation set. The dataset preserves the original 1,000-class classification task, using the same labels and image resolutions. In our setup, we focus on Gaussian noise corruption.
  • Yearbook [ginosar2015century]: This dataset contains portraits of American high school students taken over eight decades, capturing changes in visual appearance due to evolving beauty standards, cultural norms, and demographics. We follow the Wild-Time preprocessing and evaluation protocol [yao2022wild], resulting in 33,431 grayscale images (32×32 pixels) labeled with binary gender. Images from 1930–1969 are used for training, and those from 1970–2013 for testing.
  • FMoW-Time: The Functional Map of the World (FMoW) dataset [koh2021wilds] consists of 224×224 RGB satellite images categorized into 62 land-use classes. Distribution shift arises from technological and economic changes that alter land usage over time. FMoW-Time [yao2022wild] is a temporal split of FMoW-WILDS [koh2021wilds, christie2018functional], dividing 141,696 images into a training period (2002–2014) and a testing period (2015–2017).
@@ D.2 TTA Methods
  • TENT[Wang2021TentFT] updates normalization layers by minimizing test entropy.
  • SHOT [liang2020we] adapts normalization layers using information maximization and self-supervised pseudo-labeling to align target representations with a frozen source classifier.
  • SAR [niu2023towards] updates normalization layers via an entropy minimization objective. It filters out high-entropy samples and guides adaptation toward flatter minima.
  • CoTTA [wang2022continual] updates all model parameters using a student-teacher approach on augmentation averaged predictions. It also employs stochastic weight restoration to mitigate forgetting.
  • T3A [iwasawa2021test] adjusts only the final linear classifier by computing class-wise pseudo-prototypes from confident, normalized representations.
  • STAD [schirmer2024test] updates only the last linear layer by tracking the evolution of feature representations with a probabilistic state-space model.
@@ D.3 Implementation Details
We use the confseq package [confseq] by [howard2021time] to compute the conjugate-mixture empirical Bernstein confidence lower bound on the target risk. This confidence sequence framework supports tuning for an intrinsic time $t_{\text{opt}}$, which we set by default to the first 25% of the sequence length for all experiments. To implement the baseline $\hat{L}_{t}^{d}$ from jpmorgan2024sequential, we use the same loss proxy—uncertainty—as in our method.
@@ Appendix E Further Related Work
aims to assess model performance on unlabeled test data, which is often subject to distribution shift [sanyal2024accuracy, miller2021accuracy]. This is typically achieved via model uncertainty [hendrycks2016baseline, corbiere2019addressing, garg2022leveraging, lu2023predicting, guillory2021predicting] or model disagreement [jiang2021assessing, baek2022agreement, kirsch2022a, rosenfeld2023almost, lee2024aetta, nakkiran2020distributional, lee2023demystifying]. Uncertainty-based methods exploit the predictive distribution of the model— for example through the maximum class probability [hendrycks2016baseline] or the true class probability [corbiere2019addressing]—and learn a threshold to distinguish correctly from incorrectly predicted samples [garg2022leveraging, lu2023predicting, guillory2021predicting]. Disagreement-based error prediction methods leverage the theoretical equivalence between model disagreement and test error under calibration [jiang2021assessing, kirsch2022a]. However, these methods often require training multiple models—sometimes even from different architectures [baek2022agreement, jiang2021assessing]. Closest to our work [lee2024aetta, press2024entropy, kim2024test], estimate the accuracy of TTA methods based on disagreement. Notably, by exploiting theoretical results from [jiang2021assessing], lee2024aetta proposes an accuracy estimation method based on dropout disagreement. They differ from our work by (i) providing an estimator of test risk directly while we are interested in signaling a significant increase in test risk compared to the source risk; as such (ii) their method does not come with guarantees on the false alarm rate; and (iii) they require calibration (their Definition 3.3) to preserve theoretical validity of their risk estimator while we rely on separability of high and low error samples (Assumption 1).
@@
Recent work has identified several scenarios where TTA methods tend to degrade. These include adaptation under non-stationary test distributions [wang2022continual, hoang2024persistent, yuan2023robust, lim2023ttn, dobler2023robust, marsden2024universal, su2024towards, schirmer2024test], label shift [gong2022note, niu2022efficient, boudiaf2022parameter, yuan2023robust], mixed domains within a test batch [niu2023towards, lim2023ttn, marsden2024universal], small test batch sizes [niu2023towards, lim2023ttn, marsden2024universal, dobler2023robust, schirmer2024test], and adaptation in the presence of malicious samples [park2024medbn, wu2023uncovering]. Most such work on TTA robustness has focused on proposing more robust TTA methods and developing evaluation benchmarks [gong2022note, marsden2024universal, yuan2023robust, su2024towards]. liu2021ttt++ analyze failure cases of the related test-time training paradigm, which requires a self-supervised objective during training. They derive an upper bound on test risk dependent on the effectiveness of the self-supervised loss. In contrast, our approach does not require any modification to the training procedure and provides guarantees that hold regardless of the model’s original training objective. Also related to our work is research on TTA model collapse—where models degenerate to trivial solutions during adaptation [press2024entropy, niu2023towards, lee2024aetta, press2024rdumb]—and efforts to identify optimal reset mechanisms that revert the model back to its source parameters during deployment [niu2022efficient, lee2024aetta, press2024rdumb]. In contrast, we propose a general-purpose monitoring tool that provides statistical guarantees on risk control for arbitrary TTA methods. Rather than focusing on a specific mitigation strategy, our tool can inform a range of interventions—such as resetting the model to its source for continued adaptation or taking it offline entirely for retraining [hoffman2024some].